A string is stretched and fastened to two points  apart. Motion is started by displacing the string in the form  from which it is released at a time , the initial velocity is zero. To find the deflection


1
Expert's answer
2021-06-06T15:54:16-0400

"\displaystyle
\textsf{The displacement}\,\, y(x,t) \,\, \textsf{of the point of the string}\\
\textsf{at a distance}\,\, x \,\, \textsf{from the left end}\,\, 0\,\, \textsf{at time}\,\, t \\
\textsf{is given by the equation}\\

\frac{\partial^2 y}{\partial t^2} = a^2\frac{\partial^2 y}{\partial x^2}\\

\textsf{Since the ends of the string}\,\, x=0 \,\, \textsf{and}\,\, x = l \,\, \textsf{are fixed,}\\
\textsf{they do not undergo any displacement at any time.}\\

\textsf{Therefore,}
y(0,t) = 0 \,\, \textsf{for}\,\, t \geq 0 \,\, \textsf{and} \,\, y(l,t) = 0 \,\, \textsf{for}\,\, t \geq 0\\

\textsf{Since the string is released from rest initially,}
\textsf{that is, at}\,\, t=0, \,\, \textsf{the initial velocity of every point}\\
\textsf{of the string in the}\,\, y-\textsf{direction is zero.}\\

\textsf{Hence,} \frac{\partial y}{\partial t}(x,0) = 0, for 0 \leq x \leq l\\

\frac{\partial^2 y}{\partial t^2} = a^2\frac{\partial^2 y}{\partial x^2} \\

\textsf{Solving the PDE by Seperation of Variables,}\\

\textsf{Let}\,\, y = X(x)T(t)\\

\frac{\partial y}{\partial x} = X'(x)T(t)\\
\frac{\partial^2 y}{\partial x^2} = X''(x)T(t)\\
\frac{\partial y}{\partial t} = X(x)T'(t)\\
\frac{\partial^2 y}{\partial t^2} = X(x)T''(t)\\

X(x)T''(t) = a^2X''(x)T(t)\\

\frac{X''(x)}{X(x)} = \frac{T''(t)}{a^2T(t)} = K^2\\

\frac{X''(x)}{X(x)} = K^2\\

X''(x) - K^2X(x) = 0\\

\textsf{Solution of}\,\, X(x)\,\, \textsf{is}\\

X(x) = C_1\cos(Kx) + C_2\sin(Kx)\\

\frac{T''(t)}{a^2T(t)} = K^2\\

T''(t) - a^2 K^2T(t) = 0\\

T(t) = C_3\cos(aKt) + C_4\sin(aKt)\\

\textsf{Thus,}\,\,y(x,t) = \left(C_1\cos(Kx) + C_2\sin(Kx)\right)\left(C_3\cos(aKt) + C_4\sin(aKt)\right)\\

\textsf{By initial condition,}\,\,\frac{\partial y}{\partial t}(x,0) = 0\\

\frac{\partial y}{\partial t} = \left(C_1\cos(Kx) + C_2\sin(Kx)\right)\left(-C_3aK\sin(aKt) + C_4aK\cos(aKt)\right)\\

\frac{\partial y}{\partial t}(x,0) = C_4aK\left(C_1\cos(Kx) + C_2\sin(Kx)\right) = 0\\

\implies C_4 = 0.\\

\textsf{By the initial condition}\,\, y(0,t) = 0\\

y(x,t) = \left(C_1\cos(Kx) + C_2\sin(Kx)\right)\left(C_3\cos(aKt) + C_4\sin(aKt)\right) = 0\\

y(0,t) = C_1\left(C_3\cos(aKt) + C_4\sin(aKt)\right) = 0\\

\implies C_1 = 0\\

\therefore y(x,t) = C_2\sin(Kx)C_3\cos(aKt)\\

\textsf{By the initial condition}\,\, y(l,t) = 0\\

y(l,t) = C_2\sin(Kl)C_3\cos(aKt) = 0\\

\implies \sin(Kl) = 0\\

\sin(Kl) = \sin(n\pi)\hspace{1cm} \,\, \textsf{for}\,\, n \geq 1\\

K = \frac{n\pi}{l}\\

\therefore y(x,t) = C_2 C_3\sin\left(\frac{an\pi x}{l}\right)\cos\left(\frac{an\pi t}{l}\right) = \lambda\sin\left(\frac{an\pi x}{l}\right)\cos\left(\frac{an\pi t}{l}\right)\\

\textsf{The general solution is thus,}\\
y(x, t) = \sum_{n=1}^{\infty} \lambda_n \sin\left(\frac{an\pi x}{l}\right)\cos\left(\frac{an\pi t}{l}\right)\\

\textsf{Where,}\,\, \lambda_n \,\, \textsf{can be related to the Fourier}
\\ \textsf{coefficients by incorporating the initial conditions.}\\


y(x, 0) = \sum_{n=1}^{\infty} \lambda_n \sin\left(\frac{an\pi x}{l}\right)\cos\left(\frac{an\pi \times 0}{l}\right) = \sum_{n=1}^{\infty} \lambda_n \sin\left(\frac{an\pi x}{l}\right)\\

\textsf{By Fourier's Half-range sine series, we derive}\,\,\lambda_n\\

\lambda_n = \frac{2a}{l}\int_0^{\frac{l}{a}} y(x,0) \sin\left(\frac{an\pi x}{l}\right) \mathrm{d}x"